# Resisting Quantum Key Distribution Attacks Using Quantum Machine Learning

### v1

@@ III Dataset
This Section introduces a realistic QKD dataset designed to address the limitations of existing datasets, which often rely on mathematical assumptions or lack specific attack scenarios. By incorporating PNS, intercept-and-resend, Trojan-Horse, RNG, detector blinding, wavelength dependent, and combined attacks, the dataset better represents real-world threats to quantum communication. Key quantum metrics are collected to evaluate the security of the QKD process. The dataset goes through preprocessing steps, including feature selection, normalization, and augmentation, to ensure its suitability for the proposed hybrid QLSTM model. This dataset serves as a benchmark for evaluating the effectiveness of the QLSTM model in detecting QKD security threats. Existing QKD datasets (as discussed in Section I) often suffer from significant limitations, such as containing very few features, relying on mathematical assumptions, or focusing on the practical vulnerabilities of QKD systems. In addition, many studies focus on general eavesdropping threats without specifying the exact attack strategy or emphasizing the imperfections of the device. In contrast, this work explicitly incorporates PNS, Intercept-and-Resend, and Trojan-Horse attacks, which directly exploit QKD vulnerabilities at the photon transmission level. The selection of these attacks is driven by gaps identified in previous research. By addressing these threats, our dataset enables a more precise evaluation of ML-based detection mechanisms.
@@
In this work, we adopt decoy-state BB84 protocol, which enhances security by randomly transmitting additional photon (decoy) pulses with varying intensities. Alice randomly sends both signal and decoy pulses, and Bob measures them as in standard BB84. Since Eve cannot distinguish between signal and decoy pulses, she cannot selectively measure only multi-photon pulses without being detected. By analyzing detection rates for both signal and decoy states, Alice and Bob can identify and mitigate PNS attacks before finalizing the key [19].
@@
The simulation of the decoy-state BB84 protocol includes four scenarios: normal QKD operation, Intercept-and-Resend attacks, Trojan-Horse attacks, and Photon-Number-Splitting (PNS) attacks. It collects various quantum metrics, including the Quantum Bit Error Rate (QBER), bit mismatch entropy, signal and decoy detection rates, and time-based statistics, discussed in table V. The simulation is carried out to generate approximately a 256-bit secret key while reflecting practical transmission conditions. To make the dataset more realistic, different noise sources are incorporated. In particular, quantum channel losses are introduced to simulate photon attenuation over the link, photon shot noise is added to reflect the inherent fluctuations in photon detection, and depolarizing noise is applied to represent random state decoherence that naturally occurs during transmission. These noise sources vary randomly for each photon transmission, with their strength set to low, moderate, or high levels. While this does not capture every possible imperfection in a real system, it provides a closer approximation to realistic conditions compared to purely ideal simulations. Table III presents the distribution of attack scenarios in the dataset, which consists of 10,329 samples evenly spread across eight labels. In the following subsections, we describe each scenario and give details on how the effect of each condition was taken into account as we develop the dataset. Table IV summarizes the conditions we considered in each scenario.
@@
Intercept-Resend Attack Eve intercepts and measures photons before resending them to Bob. Eve randomly measures in a basis, causing errors in transmission. Increased QBER, mismatch entropy, and detection errors.
@@ III-A Attack Scenarios
In a normal DV-QKD scenario, where there is no presence of Eve, the process of exchanging the key follows the standard BB84 protocol, where Alice randomly prepares qubits in rectilinear (horizontal/vertical) or diagonal ($\pm 45^{\circ}$) bases and Bob randomly chooses a basis for measurement. After transmission, they publicly compare their chosen bases and keep only the bits where their bases match (sifting). Under these conditions, the QBER and Measurement Entropy are as close to zero as possible, indicating minimal errors and maximum predictability in the measurement outcomes. Moreover, the Signal and Decoy Detection Rates should be as close to the predefined distribution ratio of signal and decoy pulses as possible, showing that no external interference occurs. Additionally, the time-based metrics should remain stable and consistent. Any deviations from these expected values could suggest the presence of noise, losses, or an external attack on the quantum channel. Algorithm 2 illustrates the pseudo-code for the normal QKD simulation.
@@
In the Intercept-and-Resend attack scenario, the QKD process is intercepted by Eve, who measures the transmitted photons and then retransmits them to Bob, effectively impersonating Alice. This process typically introduces detectable errors, but if Eve only attacks a small fraction of signals, the errors might remain below the detection threshold. Under these conditions, the QBER and Measurement Entropy significantly increase, indicating potential errors and greater unpredictability in the measurement outcomes. Furthermore, the signal and decoy detection rates should be reduced due to Eve already measuring the signal or the decoy, which leads to possible external interference or tampering. Additionally, the time-based metrics are expected to increase because of the delay introduced by Eve’s interception. Based on these assumptions, the simulated attack is constructed as described in Algorithm 3.
@@
8:Follows Normal QKD States 9-32
@@
In the PNS attack scenario, Eve intercepts and stores a single photon from multi-photon pulses while allowing the remaining photons to reach Bob. In this attack, practical QKD systems that rely on weak coherent pulses can emit multi-photon states, which makes them vulnerable to selective photon extraction by Eve, which primarily affects the detection rates, causing a noticeable decrease. However, since Eve does not perform any measurements on the intercepted photons, the QBER and Measurement Entropy remain unchanged, as no additional errors or randomness are introduced into the quantum channel. Furthermore, the time-based metrics remain relatively stable, as the interception process does not introduce significant delays in the transmission.
@@
11:Follows Normal QKD States 9-32
Trojan Horse Attack
In the Trojan horse attack scenario, Eve injects additional photons into Alice’s system by sending bright light pulses into the QKD device on Alice’s or Bob’s side to extract internal information such as basis choice or key bits [21]. This attack primarily affects the detection rates, leading to an increase as more photons are detected on Bob’s side. However, since Eve does not perform direct measurements on the transmitted photons, the QBER and Measurement Entropy do not increase significantly, with mismatches only occurring if Bob measures an injected pulse. In addition, the time-based metrics remain stable, as the attack does not introduce notable transmission delays. Researchers in [22] indicated that Eve can adjust the photon injection strategy by varying between strong and weak pulses based on the detected reflection patterns. Based on these assumptions, the simulated attack is constructed as described in Algorithm 5.
@@
14:Follows Normal QKD States 9-32
@@
In Wavelength Dependent Trojan-Horse Attack, Eve exploits the wavelength sensitivity of QKD components by injecting bright pulses at various wavelengths, rather than only at the legitimate signal wavelength [23, 22]. This approach leverages the fact that the detection efficiency and phase encoding in optical devices can vary significantly with wavelength. By carefully selecting the injected wavelength, Eve can reduce detection efficiency, introduce phase errors, and increase her chances of eavesdropping without being detected. Unlike the classic Trojan-Horse Attack, this variant can also affect QBER and timing metrics due to wavelength-dependent delays and errors. Researchers have shown that such attacks may introduce subtle but measurable changes in the QKD statistics, potentially enabling Eve to gain information about the key while minimizing her detection risk [22, 21].
Based on these considerations, the simulated attack is constructed as described in Algorithm 6.
@@
17:Follows Normal QKD States 9-32
@@
The Random Number Generator (RNG) Attack targets the fundamental assumption that Alice and Bob’s random choices of bits and bases are truly unpredictable and uniformly distributed. If the RNGs used in the QKD protocol are biased, patterned, or otherwise predictable—due to hardware flaws or deliberate compromise—an eavesdropper (Eve) can exploit this to increase her information about the key [24, 25]. In this attack, Eve analyzes recent output patterns from the RNG and leverages statistical bias or repeated patterns to predict future choices with non-negligible probability. If Eve can correctly guess Alice’s bit and basis, she can intercept and resend the photon, introducing minimal disturbance and remaining largely undetected. The effectiveness of the attack depends on the degree of bias, the pattern predictability, and Eve’s ability to exploit them in real time. Algorithm 7 outlines the simulation process for this attack.
@@
4:  Alice generates bit and basis using compromised RNG with bias and pattern memory
@@
10:  Bob chooses basis using similarly compromised RNG
11:  Bob measures received photon and records result
@@
In Detector Blinding Attack, Eve manipulates the single-photon detectors at Bob’s side by sending bright continuous-wave light, forcing them into a linear (classical) mode where they no longer behave as true quantum detectors [26]. While blinded, Bob’s detectors lose their single-photon sensitivity and only respond to Eve’s specifically tailored trigger pulses, allowing Eve to control which detector clicks and thus gain full knowledge of the key without introducing detectable errors. The attack is typically performed in bursts (blinding duration), and Eve can inject her own signals with high efficiency while the detectors are blinded. During these periods, the detection efficiency drops, and the quantum bit error rate (QBER) may increase if the attack is not perfectly synchronized. However, if performed carefully, the attack can leave the QBER and other statistics nearly unchanged, making detection extremely challenging [26, 27]. Algorithm 8 summarizes the simulation process for this attack.
@@
The Combined Attack represents a realistic and formidable threat scenario in which an eavesdropper (Eve) employs multiple quantum hacking strategies simultaneously to maximize her chances of compromising the QKD system. In this attack, Eve leverages a combination of wavelength attacks, detector blinding, and random number generator (RNG) manipulation. By exploiting device imperfections and protocol vulnerabilities in concert, Eve can bypass optical filters with off-wavelength photons [28], force detectors into a controllable linear mode [26], and predict or bias random choices made by Alice and Bob [25]. The synergy between these attacks allows Eve to increase her information gain while minimizing detection, as the errors and losses introduced by one attack may mask or compensate for those caused by another. Algorithm 9 summarizes the simulation process for this multifaceted attack.
@@
8:  If wavelength attack active, modify photon efficiency and phase
9:  If detector blinding active, manipulate Bob’s detection probability and error rate
10:  Bob chooses basis (biased if RNG attack active) and measures photon
@@
III-B Generation
Our dataset is constructed by executing multiple iterations of the QKD simulations discussed in the previous subsections. Each iteration simulates all eight possible scenarios (normal QKD process, Intercept-and-Resend attack, PNS attack, Trojan horse attack, wavelength-dependent Trojan horse, RNG attack, detection blinding attack, and combined attack), capturing a set of QKD metrics. These include quantum bit error rate (QBER), measurement entropy, signal and decay detection rates, signal and decay loss rates, photon transmission time metrics, and time-based variation metrics. These metrics provide insights into the security of the quantum channel and help detect anomalies in the process. A total of 1292 iterations per scenario were performed, resulting in a balanced dataset of 10,329 samples. Table VI shows how the first few lines of the proposed dataset. Data preprocessing is a crucial step in the development of any ML or DL model, as it significantly affects the model’s accuracy and performance. In this context, the preprocessing framework for the proposed model contains several steps to prepare the dataset for training and evaluation. These steps include feature selection, feature transformation, dataset splitting, noise injection, and input representation.

### v2

@@ III Dataset
This Section introduces a semi-realistic QKD dataset designed to address the limitations of existing datasets, which often rely on mathematical assumptions or lack specific attack scenarios. By incorporating PNS, intercept-and-resend, Trojan-horse, RNG, detector blinding, wavelength-dependent, and combined attacks, the dataset better represents real-world threats to quantum communication, though with some limitations. Key quantum metrics are collected to evaluate the security of the QKD process. The dataset goes through preprocessing steps, including feature selection, normalization, and augmentation, to ensure its suitability for the proposed hybrid QLSTM model. This dataset serves as a benchmark for evaluating the effectiveness of the QLSTM model in detecting QKD security threats.
III-A Threat Model
We generate the dataset under a decoy-state BB84 DV-QKD adversarial framework. Eve controls the quantum channel, but has no physical access to Alice’s/Bob’s internals. She can (i) perform basis measurements and resend states (Intercept–Resend), (ii) select photon-number-dependent strategies (PNS), (iii) inject bright or off-wavelength light (Trojan-Horse variants), and (iv) drive detectors into a linear regime (Detector Blinding). Eve does not replace the quantum channel or perform loss compensation, nor does she calibrate the return flux (as in the standard THA), consistent with a bounded but capable adversary. These are side-channel and channel-level capabilities that appear through observable statistics (QBER, detection/loss rates, entropy, and timing). Consistent with [9, 22, 12, 7], we implement simplified simulation-feasible variants that preserve the operational effect of each attack while avoiding full hardware modeling. Any deviations from the strongest known forms (e.g., loss-compensation in PNS or power-calibrated THA) are documented in each scenario as modeling limitations.
III-B Simulation Setup
Existing QKD datasets (as discussed in Section I) often suffer from significant limitations, such as containing very few features, relying on mathematical assumptions, or focusing on the practical vulnerabilities of QKD systems. In addition, many studies focus on general eavesdropping threats without specifying the exact attack strategy or emphasizing device imperfections. In contrast, this work explicitly incorporates PNS, Intercept-and-Resend, and Trojan-Horse attacks, which directly exploit QKD vulnerabilities at the photon transmission level. The selection of these attacks is driven by the gaps identified in previous research. By addressing these threats, our dataset enables a more precise evaluation of ML-based detection mechanisms.
@@ III-B Simulation Setup
In this work, we adopt the decoy-state BB84 protocol, which enhances security by randomly transmitting additional photon (decoy) pulses with varying intensities. Alice randomly sends both signal and decoy pulses, and Bob measures them as in the standard BB84 protocol. Since Eve cannot distinguish between signal and decoy pulses, she cannot selectively measure only multi-photon pulses without being detected. By analyzing the detection rates for both signal and decoy states, Alice and Bob can identify and mitigate PNS attacks before finalizing the key [9].
@@
The simulation of the decoy-state BB84 protocol includes eight scenarios: Normal, Intercept-and-Resend, PNS, Trojan-Horse, Wavelength-Dependent Trojan-Horse, RNG, Detector Blinding, and Combined attacks, consistent with the threat model in Section III-A. It collects various quantum metrics, including the Quantum Bit Error Rate (QBER), bit mismatch entropy, signal and decoy detection rates, and time-based statistics, as discussed in table V. The simulation is carried out to generate approximately a 256-bit secret key while reflecting practical transmission conditions. To make the dataset more realistic, different noise sources are incorporated. In particular, quantum channel losses are introduced to simulate photon attenuation along the link, photon shot noise is added to reflect the inherent fluctuations in photon detection, and depolarizing noise is applied to model random state decoherence that naturally occurs during transmission. These noise sources vary randomly for each photon transmission, with their strength set to low, moderate, or high levels. Although this does not capture every possible imperfection in a real system, it provides a closer approximation to realistic conditions compared to purely ideal simulations. Table III summarizes the distribution of attack scenarios in the dataset, which contains 10,329 samples uniformly distributed on eight labels. The following subsections describe each scenario and explain how the impact of each condition was incorporated during dataset construction. Table IV summarizes the conditions we considered in each scenario.
@@
Intercept-Resend Attack Eve intercepts and measures photons before resending them to Bob. Eve randomly measures on a basis, causing transmission errors. Increased QBER, mismatch entropy, and detection errors.
@@ III-C Attack Scenarios
In a normal DV-QKD scenario, where there is no presence of Eve, the process of exchanging the key follows the standard BB84 protocol, where Alice randomly prepares qubits in rectilinear (horizontal/vertical) or diagonal ($\pm 45^{\circ}$) bases, and Bob randomly chooses a basis for measurement. After transmission, they publicly compare their chosen bases and keep only the bits that match (sifting). Under these conditions, the QBER and Measurement Entropy are as close to zero as possible, indicating minimal errors and maximum predictability in the measurement outcomes. Moreover, the Signal and Decoy Detection Rates should be as close as possible to the predefined distribution ratio of signal and decoy pulses, indicating that no external interference occurs. In addition, the time-based metrics should remain stable and consistent. Any deviation from these expected values could indicate noise, losses, or an external attack on the quantum channel. Algorithm 2 illustrates the pseudo-code for the normal QKD simulation.
@@
In the Intercept-and-Resend attack scenario, Eve intercepts the QKD signal, measures the transmitted photons, and then retransmits them to Bob, effectively impersonating Alice. Fundamentally, this attack exploits the direct access to the quantum channel. Eve performs a projective measurement on the in-transit qubit, effectively collapsing its quantum state. Based on her measurement outcome, she prepares and sends a new qubit to Bob. Due to the No-Cloning theorem, if Eve measures in a basis conjugate to Alice’s preparation basis, she inevitably introduces state disturbances that manifest as errors in the sifted key. This process typically introduces detectable errors, but if Eve attacks only a small fraction of signals, the errors might remain below the detection threshold. Under these conditions, the QBER and Measurement Entropy significantly increase, indicating potential errors and greater unpredictability in the measurement outcomes. Furthermore, the signal and decoy detection rates should be reduced because Eve has already measured the signal or the decoy, potentially leading to external interference or tampering. Additionally, the time-based metrics are expected to increase because of the delay introduced by Eve’s interception. Based on these assumptions, the simulated attack is constructed as described in Algorithm 3. This follows the standard BB84 analysis: full-rate Intercept–Resend would yield $\approx 25\%$ QBER, so Eve targets only a fraction of pulses to avoid alarms. We therefore expect elevated QBER/entropy and mild timing overhead, matching the metrics we log. No additional simplifications apply here.
@@
8:Follows Normal QKD States 11-32
@@
The PNS scenario follows the decoy-state model introduced by Hwang [9], in which Eve targets multi-photon pulses emitted by weak coherent sources and retains one or more photons, while allowing the remainder to reach Bob. The attack mechanism relies on the statistical nature of weak coherent laser pulses, which follow a Poisson distribution and occasionally contain more than one photon. Eve performs a Quantum Non-Demolition (QND) measurement to determine the photon number of each pulse. If she detects a multi-photon pulse, she splits off one photon to store in quantum memory and forwards the remaining photon(s) to Bob via a lossless channel. This allows her to measure the stored photon later—once the basis is announced—thereby gaining key information without inducing any bit errors. In our simulation, Eve performs photon-number–selective interception only on those multi-photon pulses; to keep the simulation computationally tractable, we do not model advanced channel-loss compensation or loss-replacement strategies that a fully resourced adversary might employ. Consequently, our traces exhibit a noticeable drop in signal/decoy detection rates, while showing little change in QBER or measurement entropy, since Eve does not measure the stored photons and therefore does not introduce additional bit errors directly. Timing metrics remain largely unaffected because the interception step introduces only a minor delay in our model. This simplified and detectable PNS variant preserves the causal link between Eve’s actions and the recorded QKD statistics (detection/loss rates, QBER, entropy), and we explicitly identify stronger covert PNS attacks that include loss compensation as out of scope for the current dataset and as limitations that will be addressed in future work.
@@
11:Follows Normal QKD States 11-32
Trojan Horse Attack (THA)
The THA simulated in this work follows the injection–reflection mechanism characterized in prior analyses [20, 11], where Eve injects bright optical pulses into Alice’s or Bob’s QKD apparatus to extract internal information such as basis settings or key bits. This side-channel attack treats the QKD device as a passive optical target. Eve launches a bright optical probe signal into the system through the quantum channel. This probe reflects off internal components (such as phase modulators), picks up the modulation state (phase or polarization) representing the current basis/bit setting, and returns it to Eve. By analyzing this back-reflected light, Eve can read out the settings without measuring the quantum signal itself. In our implementation, Eve transmits both strong and weak probe pulses without calibrating the total optical flux at Bob’s input, forming a simplified non-calibrated-power variant. In practice, a sophisticated adversary could stabilize this flux to keep detection rates constant and avoid detection; however, that level of hardware calibration is outside the simulation scope and is explicitly treated as a limitation. This non-calibrated approach intentionally introduces measurable deviations in detection and entropy metrics—representing the operational signature of an imperfectly tuned THA while preserving the correct causal relationship between Eve’s injection and the observable QKD statistics. Specifically, injected photons occasionally reach Bob’s detectors, slightly increasing the detection rate compared to normal QKD, but leaving QBER and measurement entropy largely unchanged, as Eve does not perform measurements on the transmitted quantum states. The timing metrics also remain stable, since the injection process introduces negligible latency. These simulated outcomes are consistent with the experimental observations reported in [12, 28], which demonstrated that uncalibrated THA implementations create minor but detectable statistical deviations. The omission of power calibration thus reflects a deliberate modeling simplification aimed at maintaining computational tractability while preserving the physical essence of the attack. This modeling assumption remains consistent with the bounded adversary described in Section III-A and is acknowledged as a limitation of the present study. Based on these considerations, the THA scenario is implemented as described in Algorithm 5.
Semi-Quantitative Framework for Trojan Horse Attacks. Although empirical measurements for THA parameters are not yet included in this study, we formed a quantitative framework for modeling bias and information leakage arising from injected optical probes.
Let $p_{\text{inj}}$ denote the probability that an adversary successfully injects a probing pulse within a communication slot, $\rho_{\text{ret}}$ is the effective optical return reflectivity of the internal optics, and $F_{\text{Eve}}$ is the fidelity with which the adversary infers the modulator setting from the back-reflected light. The guessing probability can characterize the resulting bias in key generation
$p_{\text{guess}}=\tfrac{1}{2}(1-p_{\text{inj}})+\big(\tfrac{1}{2}+\rho_{\text{ret}}(F_{\text{Eve}}-\tfrac{1}{2})\big)p_{\text{inj}},$
which defines the conditional min-entropy per bit as
$H_{\min}(K|E)=-\log_{2}p_{\text{guess}},$
and the corresponding entropy loss $\delta H_{\min}=1-H_{\min}(K|E)$. The expected information leakage for an $n$-bit raw key can then be upper-bounded by
$L_{\max}=n(1-H_{\min}).$
These expressions provide an analytical framework to evaluate how probe-injection frequency, component reflectivity, and adversarial inference fidelity jointly influence the effective security of QKD systems under THA scenarios. Future simulation or experimental data can be used directly to estimate entropy reduction and leakage levels.
@@
14:Follows Normal QKD States 11-32
@@
The wavelength-dependent Trojan-Horse Attack simulated in this work targets the spectral sensitivity of QKD components rather than relying on reflected power at the legitimate wavelength, as in the standard THA. In this variant, Eve injects optical pulses at slightly offset wavelengths to exploit wavelength-dependent variations in detector efficiency and phase modulation [7, 12]. This attack method takes advantage of the chromatic dispersion and wavelength-dependent coupling ratios of optical components. Optical devices such as beam splitters and modulators operate nominally at a specific design wavelength; shifting the probe wavelength allows Eve to alter splitting ratios, bypass spectral filters, or induce phase shifts that differ from the operational norm. This enables her to extract information or manipulate the system by finding spectral holes in the hardware defenses. For each injected wavelength $\lambda_{\text{attack}}$, the simulation computes the wavelength offset $\Delta\lambda=|\lambda_{\text{attack}}-\lambda_{\text{legit}}|$ and updates the detector response through efficiency $\eta(\Delta\lambda)$ and phase $\phi(\Delta\lambda)$. These wavelength-induced perturbations can reduce detection efficiency, introduce phase errors, and slightly increase photon arrival-time dispersion—effects that manifest directly in QBER, detection/loss rates, and timing metrics. While both THA variants share the same injection–reflection principle, their physical mechanisms and statistical footprints differ: the standard THA manipulates total optical flux, whereas the wavelength-dependent variant alters spectral response, producing distinct phase and timing signatures. This distinction allows the hybrid QLSTM model to recognize spectral leakage patterns without conflating them with flux-based deviations. To maintain computational feasibility, the simulation adopts a parametric model that captures the causal link between injected wavelength shifts and measurable QKD statistics, without implementing full optical modeling such as modal coupling or detector spectral calibration. More advanced, hardware-calibrated implementations of this attack are explicitly considered beyond the scope of this dataset and are listed as limitations of the present study.
@@
17:Follows Normal QKD States 11-32
@@
The Random Number Generator (RNG) Attack targets the fundamental assumption that Alice and Bob’s random choices of bits and bases are truly unpredictable and uniformly distributed. If the RNGs used in the QKD protocol are biased, patterned, or otherwise predictable—due to hardware flaws or deliberate compromise—an eavesdropper (Eve) can exploit this to increase her information about the key [13, 6]. The attack capitalizes on non-random correlations or biases in the RNG output stream. Eve collects a history of public basis announcements and, if available, compromised key bits, to train a predictive model (e.g., using statistical analysis or machine learning). By forecasting the basis choices of Alice and Bob with high probability, Eve can align her intercept-resend measurements with the predicted basis, drastically reducing the error rate (QBER) that typically reveals her presence. In this attack, Eve analyzes recent output patterns from the RNG and leverages statistical bias or repeated patterns to predict future choices with non-negligible probability. If Eve can correctly guess Alice’s bit and basis, she can intercept and resend the photon, introducing minimal disturbance and remaining largely undetected. The effectiveness of the attack depends on the degree of bias, the predictability of the pattern, and Eve’s ability to exploit them in real time. The Algorithm 7 describes the simulation process for this attack.
@@
4:  Alice generates a bit and basis using a compromised RNG with bias and pattern memory
@@
10:  Bob chooses a basis using a similarly compromised RNG
11:  Bob measures received photon and records the result
@@
In Detector Blinding Attack, Eve manipulates the single-photon detectors on Bob’s side by sending bright continuous-wave light, forcing them into a linear (classical) mode where they no longer behave as true quantum detectors [22]. The core principle is to overwhelm the Avalanche Photodiodes (APDs) with high-intensity illumination, preventing them from operating in Geiger mode, where they are sensitive to single photons. Once blinded, the detectors are essentially turned off until a pulse with energy exceeding a high discrimination threshold arrives. Eve can then send bright ’trigger’ pulses precisely when she wants a detection to occur, effectively taking full control of Bob’s measurement outcomes without his knowledge. While blinded, Bob’s detectors lose their single-photon sensitivity and only respond to Eve’s specifically tailored trigger pulses, allowing Eve to control which detector clicks, and thus gain full knowledge of the key without introducing detectable errors. The attack is typically performed in bursts (blinding duration), and Eve can inject her own signals with high efficiency while the detectors are blinded. During these periods, the detection efficiency drops, and the quantum bit error rate (QBER) may increase if the attack is not perfectly synchronized. However, if performed carefully, the attack can leave the QBER and other statistics nearly unchanged, making detection extremely challenging [22, 23]. The Algorithm 8 summarizes the simulation process for this attack. We follow [22]: bright CW light forces a linear regime ($\eta_{\text{blinded}}\ll\eta_{\text{normal}}$) and Eve injects triggers. In ideal synchronization, QBER may remain flat;however, occasional desynchronization yields slight QBER increases and reduced efficiency, both of which are observable in our features.
@@
The Combined Attack represents a realistic and formidable threat scenario in which an eavesdropper (Eve) simultaneously employs multiple quantum-hacking strategies to maximize her chances of compromising the QKD system. In this attack, Eve leverages a combination of wavelength attacks, detector blinding, and manipulation of a random number generator (RNG). This compound strategy operates by layering distinct attack vectors to cover each other’s weaknesses. For instance, the predictive power of an RNG attack can reduce the error rate inherent in intercept-resend maneuvers, while detector blinding enables Eve to control the exact timing and success of her injected pulses, masking the statistical anomalies (like loss rates) that usually accompany other active attacks. By exploiting device imperfections and protocol vulnerabilities in concert, Eve can bypass optical filters with off-wavelength photons [32], force detectors into a controllable linear mode [22], and predict or bias random choices made by Alice and Bob [6]. The synergy between these attacks allows Eve to increase her information gain while minimizing detection, as the errors and losses introduced by one attack may mask or compensate for those caused by another. The Algorithm 9 summarizes the simulation process for this multifaceted attack.
@@
8:  If wavelength attack is active, modify photon efficiency and phase
9:  If detector blinding is active, manipulate Bob’s detection probability and error rate
10:  Bob chooses a basis (biased if RNG attack is active) and measures a photon
@@
15:Follows Normal QKD States 11-32
III-D Generation
Our dataset is constructed by executing multiple iterations of the QKD simulations discussed in the previous subsections. Each iteration simulates all eight possible scenarios (normal QKD process, Intercept-and-Resend attack, PNS attack, Trojan horse attack, wavelength-dependent Trojan horse, RNG attack, detection blinding attack, and combined attack) and captures a set of QKD metrics. These include quantum bit error rate (QBER), measurement entropy, signal and decoy detection rates, signal and decoy loss rates, photon transmission time metrics, and time-based variation metrics. These metrics provide insights into the security of the quantum channel and help detect anomalies in the process. A total of 1292 iterations were performed per scenario, resulting in a balanced dataset of 10,329 samples. Table VI shows the first lines of the proposed dataset. Data preprocessing is a crucial step in developing any ML or DL model, as it significantly affects the model’s accuracy and performance. In this context, the preprocessing framework for the proposed model contains several steps to prepare the dataset for training and evaluation. These steps include feature selection, feature transformation, dataset splitting, noise injection, and input representation.